OpenStudy (anonymous):

\[x^{2}+3x+5=0\]and\[ax^{2}+bx+c=0\]have a common root and \[a,b,c \epsilon N\]then the minimum value of a+b+c is ???

OpenStudy (anonymous):

i am getting 9

OpenStudy (anonymous):

is it correct @mukushla ?

OpenStudy (anonymous):

show ur work

OpenStudy (anonymous):

i found that both the roots are common

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

eq. 1 will have 2 complex roots so both the roots will be equal so 1/a = 3/b = 5/c

OpenStudy (anonymous):

i mean to say both the roots will be common

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

minimum a+b+c=1+3+5=9